# Almost sure convergence rates of stochastic gradient methods under gradient domination

### v1

@@ A.4 Auxiliary Convergence Theorems
for $n\in\mathbb{N}$. Using Theorem A.1 we observe that there exists $Z_{\infty}$ almost surely finite such that $Z_{n}=r_{n}Y_{n}\to Z_{\infty}$ almost surely as $n\to\infty$. Moreover, we obtain that
@@
Appendix B Proof of Lemma 3.1
Proof.
@@
such that the claim follows by [27, Lem. 1].
@@
which was also applied and proved in [27, Lem. 1]. The aim is to apply the Robbins-Siegmund corollary, Lemma A.2, in order to derive the almost sure convergence rate. Let $21\leq q<$ 2 be arbitrary for now. The key step of the proof is the following computation
@@
Next, we apply the elementary inequality (9) and choose $q$ such that $\frac{1}{2}<\theta\leq\frac{1}{q}\leq 1$. Moreover by the choice of $\gamma_{n}$, there exists some $c_{4}>0$ such that $\tilde{c}_{1}\gamma_{n}^{q}\geq\frac{c_{4}}{t^{q\theta}}$ for all $n\geq N$. It follows that for all $n\geq N$
@@
The first condition (15) is obviously satisfied, since we assume $\theta\leq\frac{1}{q}$. For the second condition (16) we may choose $\theta>1-\frac{\eta}{2}$ such that $11-\eta-2\theta<-$ 1. The third condition (17) gives $11-\eta-\frac{\theta(2\beta q-1)}{2\beta-1}<-$ 1 which leads to the condition $\theta>\frac{(2-\eta)(2\beta-1)}{2\beta q-1}$. Hence, all together we obtain the sufficient condition
@@
Appendix C Proof of Theorem 4.1
Proof.
Recall, in Section 2 we derived Equation (5),
$\mathbb{E}[f(X_{n+1})-f^{\ast}\mid\mathcal{F}_{n}]$
which will be the basis of the proof of Theorem 4.1.
We treat again both cases for $\beta=\frac{1}{2}$ and $\beta\in(\frac{1}{2},1]$ separately:
$\beta=\frac{1}{2}$: In this case, Equation (5) results in the super-martingale inequality
$\beta\in(\frac{1}{2},1]$: In this case, Equation (5) results in the super-martingale inequality
For claim $(ii)$ we firstly proceed as in the proof of Lemma 3.1. Therefore, one can choose the auxiliary parameter $1<q\leq\frac{1}{\theta}$ and find constants $c_{4},c_{3},\tilde{N}_{1}>0$ such that for all $n\geq\tilde{N}_{1}$ by Equation(14) we have
Next, we take the expectation to obtain
for all $n\geq\tilde{N}_{1}$, implying that $w_{n}=\mathbb{E}[n^{1-\eta}Y_{n}]\to 0$ as $n\to\infty$ by Lemma A.3. Note that we have chosen $\theta,\eta$ and $q$ as in Lemma 3.1, such that $\sum_{n}\frac{1}{n^{q\theta}}=\infty$, $\sum_{n}(n+1)^{1-\eta}\gamma_{n}^{\frac{2\beta q-1}{2\beta-1}}<\infty$, and $\sum_{n}(n+1)^{1-\eta}\gamma_{n}^{2}<\infty$ (see (15), (16) and (17)). Therefore, the assumptions of Lemma A.3 are met.
Appendix D Proof of Theorem 4.2
Recall the definition of the iteration scheme (SHB). Using the following definitions
$Z_{n}=X_{n}+\frac{\nu}{1-\nu}W_{n},\quad W_{n}=X_{n}-X_{n-1},$ (18)
one can derive the iterative evolution
$W_{n+1}=\nu W_{n}-\gamma_{n}V(X_{n})$ (19)
$Z_{n+1}=Z_{n}-\frac{\gamma_{n}}{1-\nu}V(X_{n}).$ (20)
We will utilize these auxiliary variables in the proof.
Proof.
The proof begins as in the proof of Theorem 2 in [27]. Using only $L$-smoothness and assumption (ABC), they show that for any $c_{3}\in(0,\frac{1}{1-\nu})$, $\lambda\in(\nu,1)$ there exist constants $c_{1},c_{2},c_{4}>0$ such that choosing the step size $\gamma_{n}\sim\frac{1}{n^{\theta}}$, for some $\theta\in(\frac{1}{2},1)$ results in [27, Equation (21)]
$\beta=\frac{1}{2}$: Instead of $\mu$-strong convexity we use the gradient domination inequality $\lVert\nabla f(x)\rVert^{2}\geq c(f^{\ast}-f(x))$, as the same inequality is implied by strong convexity using $c=\mu$. Then, Claim (i) follows using the same proof as [27, Thm. 2b)]. Note that the inequality
$\frac{1}{2L}\lVert\nabla f(x)\rVert^{2}\leq f(x)-f^{\ast},$ (23)
used in the last step only requires the $L$-smoothness assumption [38, Sec. 1.2.3].
For Claim (ii) we consider Equation (22) which simplifies for $\beta=\frac{1}{2}$ to
Let $c_{5}=\min\{\tilde{c}_{1},\tilde{c}_{2}\}$, multiply by $(n+1)^{1-\eta}$ on both sides and use Equation (9) to obtain for $n\geq N$
$\leq\left(1-\tilde{c}_{5}\gamma_{n}\right)\mathbb{E}\left[n^{1-\eta}\left(Q_{n}+\lVert W_{n}\rVert^{2}\right)\right]+c_{4}\gamma_{n}^{2}(n+1)^{1-\eta}.$
To finish the proof, one can derive
$f(X_{n})-f^{\ast}\leq Q_{n}+\frac{1}{2}\lVert\nabla f(X_{n})\rVert^{2}+\frac{\nu^{2}+L\nu^{2}}{1(1-\nu)^{2}}\lVert W_{n}\rVert^{2}$ (24)
see [27, Equation (19)] for more details. Using inequality (23), we get almost surely
$\left(1-\frac{1}{4L}\right)f(X_{n})-f^{\ast}\leq Q_{n}+\frac{\nu^{2}+L\nu^{2}}{1(1-\nu)^{2}}\lVert W_{n}\rVert^{2}.$ (25)
$\beta\in(\frac{1}{2},1]$: For Claim (i), note that in Equation (22) $\lambda<1$, such that
$\mathbb{E}[Q_{n+1}+\lVert W_{n+1}\rVert^{2}\mid\mathcal{F}_{n}]$
By Lemma 3.1 we obtain that $Q_{n}+\lVert W_{n}\rVert^{2}=f(Z_{n})-f^{\ast}+\lVert W_{n}\rVert^{2}\in o\left(\frac{1}{n^{1-\eta}}\right)$ for all $\eta\in\left(\max\{2-2\theta,\frac{\theta+2\beta-2}{2\beta-1}\},1\right)$. We apply the inequality in (25) to conclude that also $f(X_{n})-f^{\ast}\in o\left(\frac{1}{n^{1-\eta}}\right)$ for all $\eta\in\left(\max\{2-2\theta,\frac{\theta+2\beta-2}{2\beta-1}\},1\right)$. This proves Claim (i).
For Claim (ii), we again use the $q$-trick from Lemma 3.1 in Equation (22). For $21<q<\frac{1}{\theta}<$ 2 we have that
$\mathbb{E}[Q_{n+1}+\lVert W_{n+1}\rVert^{2}\mid\mathcal{F}_{n}]$
Now with (12) in Lemma 3.1 there exists $\tilde{c}_{3}\geq 0$ such that
From the proof of Lemma 3.1, we choose the auxiliary parameter $q$ such that $\sum_{n}(n+1)^{1-\eta}\left(\gamma_{n}^{\frac{2\beta q-1}{2\beta-1}}+\gamma_{n}^{2}\right)<\infty$ (see (16) and (17)). By applying again Lemma A.3 we obtain $\mathbb{E}[n^{1-\eta}(Q_{n}+\lVert W_{n}\rVert^{2})]\to 0$, i.e. $\mathbb{E}[Q_{n}+\lVert W_{n}\rVert^{2}]\in o(\frac{1}{n^{1-\eta}})$. Finally, Claim (ii) follows again by Equation (25). ∎
Appendix E Numerical experiment - Toy example
Figure 1: Pathwise error $(f_{p}(X_{n}))_{n=1,\dots,N}$ of SGD and SHB for various choices of $\beta\in\{0.5,0.67,0.83,0.92\}$. For each setting we have simulated $100$ runs of length , $N=10^{5}$. The black dash-dotted line corresponds to the theoretical rate $n^{-\frac{1}{4\beta-1}}$.
We have implemented the same toy example similar to [12] to test our theoretical findings. In our implementation, we consider both SGD and SHB applied to the objective function $f_{p}(x)=|x|^{p}$, where $x\in\mathbb{R}$, for various choices of $p\geq 2$. It is straightforward to verify that $f_{p}$ satisfies the global gradient domination with parameter $\beta(p)=\frac{p-1}{p}$. It is noteworthy that for $p=2$, the $f_{p}$ obviously satisfies the PL condition with $\beta=\frac{1}{2}$, whereas for increasing $p\to\infty$, we move towards $\beta(p)\to 1$. We have used the step size schedule $\Theta(n^{-\frac{2\beta(p)}{4\beta(p)-1}})$ discussed in Table 1 and observed the almost sure convergence rates $n^{-\frac{1}{4\beta(p)-1}}$ as suggested by Theorem 4.1 and Theorem 4.2. Note that our derived rates are arbitrarily close to the sharp upper bound known in expectation [12].
Details of the implementation:
Appendix F Proof of Section 5
F.1 Proof of Theorem 5.1
Suppose that the assumptions of Theorem 5.1 hold throughout this section.
@@ Appendix F Proof of Section 5
In contrast to the global gradient domination analysis we may assume w.l.o.g. the uniform second moment bounds, i.e. $A=B=0$, instead of the more general (ABC) condition. Choosing $A,B>0$ would imply the bounded variance assumption of the gradient estimator. Note therefore, that the first term $A(f(x)-f(x^{\ast}))$ and the second term $B\lVert\nabla f(x)\rVert^{2}$ are both locally bounded by the local Lipschitz assumptions on $f$ and $\nabla f$.
@@
The outline of the proof is structured as follows:
  • First, we unify the gradient domination property around the set of local minima $\mathcal{X}^{\ast}$ and obtain a radius $r$ such that the unified gradient domination property is fulfilled in all open balls with radius $r$ around $x^{\ast}\in\mathcal{X}^{\ast}$ (Lemma F.1).
  • Based on this we construct sets $\mathcal{U}$, $\mathcal{U}_{1}\subseteq\mathbb{R}^{d}$ and the events $\Omega_{n}\in\Omega$ (see (26), (27) and (28)), such that $\Omega_{\mathcal{U}}=\bigcap_{n}\Omega_{n}$ occurs with high probability. To be precise, $\mathcal{U}_{1}$ and $\mathcal{U}$ are neighborhoods of $\mathcal{X}^{\ast}$ constructed such that the gradient domination property holds within this region, and when starting in $\mathcal{U}_{1}$ the gradient trajectory does remain in $\mathcal{U}$ for all gradient steps with high probability. Then, $\Omega_{n}$ describes the event that $X_{k}\in\mathcal{U}$ for all $k\leq n$.
@@
  • In order to show $\mathbb{P}(\Omega_{n})\geq 1-\delta$ we construct set $C_{n}$ and $E_{n}$ defined in (29) and (36) such that $E_{n}\cap C_{n}\subset\Omega_{n+1}$ (Lemma F.6) while Lemma F.5 is used to prove this claim.
  • The sets $E_{n}$ are such that $f(X_{n})$ remains close to $f^{\ast}$. We exploit the unified gradient domination property to construct the sets $E_{n}$ (Lemma F.4) and derive a recursive inequality in Lemma F.6 c) to prove that this event occurs with high probability (Lemma F.7).
  • The sets $C_{n}$ are such that $X_{n+1}$ remains close to $X_{n}$ and we exploit the finite variance assumption to show that these events occur with high probability (Lemma F.8).
@@
Remark F.2.
It is noteworthy that the unified gradient domination property obtained in the previous Lemma does not require an absolute value, as $f(x)\geq l$ for all $x\in\bigcup_{x^{\ast}\in\mathcal{X}^{\ast}}\tilde{\mathcal{B}}_{r}(x^{\ast})$. This is crucial to obtain the recursive inequalities in Lemma F.4 and we will exploit this also in the proof of Theorem 5.1 to obtain the convergence rates.
@@
In order to prove Theorem 5.1 we will show that $\Omega_{n}$ has probability at least $1-\delta$ for all $n\in\mathbb{N}$. To do this, we construct another sequence of events $(\hat{E}_{n})$ with $\hat{E}_{n}\subset\Omega_{n}$ which occurs with probability at least $1-\delta$ for any $n\in\mathbb{N}$.
[…]
Lemma F.4.
@@
$\tilde{D}_{n+1}$ $\leq(1-\gamma_{n}c^{2})\tilde{D}_{n}+\gamma_{n}\xi_{n}\mathbf{1}_{\Omega_{n}}+\frac{L\gamma_{n}^{2}}{2}\mathbf{1}_{\Omega_{n}}\lVert V_{n+1}(X_{n})\rVert^{2},$ (30)
$\leq\tilde{D}_{1}\prod_{k=1}^{n}(1-\gamma_{k}c^{2})+\sum_{k=1}^{n}\gamma_{k}\xi_{k}\mathbf{1}_{\Omega_{k}}+\frac{L}{2}\sum_{k=1}^{n}\gamma_{k}^{2}\lVert V_{k+1}(X_{k})\rVert^{2}\mathbf{1}_{\Omega_{k}}.$ (31)
@@
$\tilde{D}_{n+1}$ $\leq(1-\gamma_{n}^{q}c^{2})\tilde{D}_{n}+(2\beta)^{-\frac{1}{2\beta-1}}(1-\frac{1}{2\beta}){c^{2}}\gamma_{n}^{\frac{2\beta q-1}{2\beta-1}}+\gamma_{n}\xi_{n}\mathbf{1}_{\Omega_{n}}+\frac{L\gamma_{n}^{2}}{2}\lVert V_{n+1}(X_{n})\rVert^{2}$ (32)
$\leq\tilde{D}_{1}\prod_{k=1}^{n}(1-\gamma_{k}^{q}c^{2})+\tilde{c}\sum_{k=1}^{n}\gamma_{k}^{\frac{2\beta q-1}{2\beta-1}}+\sum_{k=1}^{n}\gamma_{k}\xi_{k}\mathbf{1}_{\Omega_{k}}+\frac{L}{2}\sum_{k=1}^{n}\gamma_{k}^{2}\lVert V_{k+1}(X_{k})\rVert^{2},$ (33)
@@
for $Z(X_{n},\zeta_{n+1})$ from 2.4 and $\xi_{n+1}=-\langle\nabla f(X_{n}),Z(X_{n},\zeta_{n+1})\rangle$.
@@
$\begin{split}\tilde{D}_{n+1}&\leq D_{n+1}\mathbf{1}_{\Omega_{n}}\\ &\leq D_{n}\mathbf{1}_{\Omega_{n}}-\gamma_{n}\mathbf{1}_{\Omega_{n}}\lVert\nabla f(X_{n})\rVert^{2}+\gamma_{n}\mathbf{1}_{\Omega_{n}}\xi_{n+1}+\frac{L\gamma_{n}^{2}}{2}\mathbf{1}_{\Omega_{n}}\lVert V_{n+1}(X_{n})\rVert^{2}\\ &\leq D_{n}\mathbf{1}_{\Omega_{n}}-\gamma_{n}c^{2}(f(X_{n})-l)\mathbf{1}_{\Omega_{n}}+\gamma_{n}\mathbf{1}_{\Omega_{n}}\xi_{n+1}+\frac{L\gamma_{n}^{2}}{2}\mathbf{1}_{\Omega_{n}}\lVert V_{n+1}(X_{n})\rVert^{2}\\ &=(1-\gamma_{n}c^{2})\tilde{D}_{n}+\gamma_{n}\xi_{n+1}\mathbf{1}_{\Omega_{n}}+\frac{L\gamma_{n}^{2}}{2}\mathbf{1}_{\Omega_{n}}\lVert V_{n+1}(X_{n})\rVert^{2},\\ &\leq\tilde{D}_{1}\prod_{k=1}^{n}(1-\gamma_{k}c^{2})+\sum_{k=1}^{n}\gamma_{k}\xi_{k+1}\mathbf{1}_{\Omega_{k}}+\frac{L}{2}\sum_{k=1}^{n}\gamma_{k}^{2}\lVert V_{k+1}(X_{k})\rVert^{2}\mathbf{1}_{\Omega_{k}},\end{split}$ (34)

### v3

@@ Appendix A Auxiliary Convergence Theorems
for $n\in\mathbb{N}$. Using Lemma A.1 we observe that there exists $Z_{\infty}$ almost surely finite such that $Z_{n}=r_{n}Y_{n}\to Z_{\infty}$ almost surely as $n\to\infty$. Moreover, we obtain that
@@
Appendix B Numerical experiment - Toy example
In the following numerical experiment, we aim to verify our theoretical finding. We have implemented the same toy example similar to [11] to test our theoretical findings. In our implementation, we consider both SGD and SHB applied to the objective function $f_{p}(x)=|x|^{p}$, where $x\in\mathbb{R}$, for various choices of $p\geq 2$. It is straightforward to verify that $f_{p}$ satisfies the global gradient domination with parameter $\beta(p)=\frac{p-1}{p}$. It is noteworthy that for $p=2$, the $f_{p}$ obviously satisfies the PL condition with $\beta=\frac{1}{2}$, whereas for increasing $p\to\infty$, we move towards $\beta(p)\to 1$. We have used the step size schedule $\Theta(n^{-\frac{2\beta(p)}{4\beta(p)-1}})$ discussed in Table 1 and observed the almost sure convergence rates $n^{-\frac{1}{4\beta(p)-1}}$ as suggested by Theorem 4.1 and Theorem 4.2. Note that our derived rates are arbitrarily close to the sharp upper bound known in expectation [11].
Figure 1: Pathwise error $(f_{p}(X_{n}))_{n=1,\dots,N}$ of SGD and SHB for various choices of $\beta\in\{0.5,0.67,0.83,0.92\}$. For each setting we have simulated $100$ runs of length , $N=10^{5}$. The bold lines correspond to the average error of SGD (red) and SHB (blue), and the black dash-dotted line corresponds to the theoretical rate $n^{-\frac{1}{4\beta-1}}$.
Details of the implementation:
Appendix C Proof of Lemma 3.1
In the following section, we present the proof of our super-martingale result.
Proof of Lemma 3.1.
@@ Appendix B Numerical experiment - Toy example
such that the claim follows by Liu and Yuan [28, Lem. 1].
@@
which was also applied and proved in Liu and Yuan [28, Lem. 1]. The aim is to apply the Robbins-Siegmund implication, Lemma A.2, in order to derive the almost sure convergence rate. Let $21\leq q<$ 2 be arbitrary for now. The key step of the proof is the following computation
@@
Next, we apply the elementary inequality Equation 17 and choose $q$ such that $\frac{1}{2}<\theta\leq\frac{1}{q}\leq 1$. Moreover by the choice of $\gamma_{n}$, there exists some $c_{4}>0$ such that $\tilde{c}_{1}\gamma_{n}^{q}\geq\frac{c_{4}}{t^{q\theta}}$ for all $n\geq N$. It follows that for all $n\geq N$
@@
The first condition Equation 23 is obviously satisfied, since we assume $\theta\leq\frac{1}{q}$. For the second condition Equation 24 we may choose $\theta>1-\frac{\eta}{2}$ such that $11-\eta-2\theta<-$ 1. The third condition Equation 25 gives $11-\eta-\frac{\theta(2\beta q-1)}{2\beta-1}<-$ 1 which leads to the condition $\theta>\frac{(2-\eta)(2\beta-1)}{2\beta q-1}$. Hence, all together we obtain the sufficient condition
@@
Appendix D Proofs of Section 5
In the following section, we present the proofs of Section 5.
D.1 Proof of Theorem 5.1
@@ Appendix D Proofs of Section 5
Suppose that the assumptions of Theorem 5.1 hold throughout this section. In contrast to the global gradient domination analysis we may assume w.l.o.g. the uniform second moment bounds, i.e. $A=B=0$, instead of the more general Equation ABC condition. Choosing $A,B>0$ would imply the bounded variance assumption of the gradient estimator. Note therefore, that the first term $A(f(x)-f(x^{\ast}))$ and the second term $B\lVert\nabla f(x)\rVert^{2}$ are both locally bounded by the local Lipschitz assumptions on $f$ and $\nabla f$.
@@
Recall, the outline of the proof is structured as follows:
  • First, we unify the gradient domination property around the set of local minima $\mathcal{X}^{\ast}$ and obtain a radius $r$ such that the unified gradient domination property is fulfilled in all open balls with radius $r$ around $x^{\ast}\in\mathcal{X}^{\ast}$ (Lemma D.1).
  • Based on this we construct sets $\mathcal{U}$, $\mathcal{U}_{1}\subseteq\mathbb{R}^{d}$ and the events $\Omega_{n}\in\Omega$ (see Equation 26, Equation 27 and Equation 28), such that $\Omega_{\mathcal{U}}=\bigcap_{n}\Omega_{n}$ occurs with high probability. To be precise, $\mathcal{U}_{1}$ and $\mathcal{U}$ are neighborhoods of $\mathcal{X}^{\ast}$ constructed such that the gradient domination property holds within this region, and when starting in $\mathcal{U}_{1}$ the gradient trajectory does remain in $\mathcal{U}$ for all gradient steps with high probability. Then, $\Omega_{n}$ describes the event that $X_{k}\in\mathcal{U}$ for all $k\leq n$.
@@
  • In order to show $\mathbb{P}(\Omega_{n})\geq 1-\delta$ we construct set $C_{n}$ and $E_{n}$ defined in Equation 29 and Equation 34 such that $E_{n}\cap C_{n}\subset\Omega_{n+1}$ (Lemma D.6) while Lemma D.5 is used to prove this claim.
  • The sets $E_{n}$ are such that $f(X_{n})$ remains close to $f^{\ast}$. We exploit the unified gradient domination property to construct the sets $E_{n}$ (Lemma D.4) and derive a recursive inequality in Lemma D.6 c) to prove that this event occurs with high probability (Lemma D.7).
  • The sets $C_{n}$ are such that $X_{n+1}$ remains close to $X_{n}$ and we exploit the finite variance assumption to show that these events occur with high probability (Lemma D.8).
@@
Remark D.2.
It is noteworthy that the unified gradient domination property obtained in the previous Lemma does not require an absolute value, as $f(x)\geq l$ for all $x\in\bigcup_{x^{\ast}\in\mathcal{X}^{\ast}}\tilde{\mathcal{B}}_{r}(x^{\ast})$. This is crucial to obtain the recursive inequalities in Lemma D.4 and we will exploit this also in the proof of Theorem 5.1 to obtain the convergence rates.
@@
In order to prove Theorem 5.1 we will show that $\Omega_{n}$ has probability at least $1-\delta$ for all $n\in\mathbb{N}$. To do this, we construct another sequence of events $(\hat{E}_{n})$ with $\hat{E}_{n}\subset\Omega_{n}$ which occur with probability at least $1-\delta$ for any $n\in\mathbb{N}$.
Therefore, we fix the notation $D_{n}:=f(X_{n})-l$, $\xi_{n+1}:=-\langle\nabla f(X_{n}),Z(X_{n},\zeta_{n+1})\rangle$ and $\mathbf{1}_{\mathcal{A}}$ denotes the indicator function for a measurable set $\mathcal{A}$ in $(\Omega,\mathcal{F},\mathbb{P})$, i.e. $\mathbf{1}_{\mathcal{A}}(\omega)=1$ if $\omega\in\mathcal{A}$ and $\mathbf{1}_{\mathcal{A}}(\omega)=0$ if $\omega\notin\mathcal{A}$. We prove the following (recursive) inequalities.
Lemma D.4.
@@
$\begin{split}{D}_{n+1}\mathbf{1}_{\Omega_{n}}&\leq(1-\gamma_{n}c^{2})D_{n}\mathbf{1}_{\Omega_{n}}+\gamma_{n}\xi_{n+1}\mathbf{1}_{\Omega_{n}}+\frac{L\gamma_{n}^{2}}{2}\mathbf{1}_{\Omega_{n}}\lVert V_{n+1}(X_{n})\rVert^{2},\\ &\leq{D}_{1}\prod_{k=1}^{n}(1-\gamma_{k}c^{2})\mathbf{1}_{\Omega_{n}}+\sum_{k=1}^{n}\left(\prod_{j=k}^{n}(1-\gamma_{j}c^{2})\right)\gamma_{k}\xi_{k+1}\mathbf{1}_{\Omega_{n}}\\ &\quad+\frac{L}{2}\sum_{k=1}^{n}\gamma_{k}^{2}\lVert V_{k+1}(X_{k})\rVert^{2}\mathbf{1}_{\Omega_{n}}.\end{split}$ (30)
@@
$\begin{split}{D}_{n+1}\mathbf{1}_{\Omega_{n}}&\leq(1-\gamma_{n}^{q}c^{2}){D}_{n}\mathbf{1}_{\Omega_{n}}+(2\beta)^{-\frac{1}{2\beta-1}}(1-\frac{1}{2\beta}){c^{2}}\gamma_{n}^{\frac{2\beta q-1}{2\beta-1}}+\gamma_{n}\xi_{n+1}\mathbf{1}_{\Omega_{n}}+\frac{L\gamma_{n}^{2}}{2}\lVert V_{n+1}(X_{n})\rVert^{2}\mathbf{1}_{\Omega_{n}}\\ &\leq{D}_{1}\prod_{k=1}^{n}(1-\gamma_{k}^{q}c^{2})+\tilde{c}\sum_{k=1}^{n}\gamma_{k}^{\frac{2\beta q-1}{2\beta-1}}+\sum_{k=1}^{n}\left(\prod_{j=k}^{n}(1-\gamma_{j}^{q}c^{2})\right)\gamma_{k}\xi_{k+1}\mathbf{1}_{\Omega_{n}}\\ &\quad+\frac{L}{2}\sum_{k=1}^{n}\gamma_{k}^{2}\lVert V_{k+1}(X_{k})\rVert^{2}\mathbf{1}_{\Omega_{n}},\end{split}$ (31)
@@
for $Z(X_{n},\zeta_{n+1})$ from Assumption 2.4 and $\xi_{n+1}=-\langle\nabla f(X_{n}),Z(X_{n},\zeta_{n+1})\rangle$.
@@
$\begin{split}D_{n+1}\mathbf{1}_{\Omega_{n}}&\leq D_{n}\mathbf{1}_{\Omega_{n}}-\gamma_{n}\mathbf{1}_{\Omega_{n}}\lVert\nabla f(X_{n})\rVert^{2}+\gamma_{n}\mathbf{1}_{\Omega_{n}}\xi_{n+1}+\frac{L\gamma_{n}^{2}}{2}\mathbf{1}_{\Omega_{n}}\lVert V_{n+1}(X_{n})\rVert^{2}\\ &\leq D_{n}\mathbf{1}_{\Omega_{n}}-\gamma_{n}c^{2}(f(X_{n})-l)\mathbf{1}_{\Omega_{n}}+\gamma_{n}\mathbf{1}_{\Omega_{n}}\xi_{n+1}+\frac{L\gamma_{n}^{2}}{2}\mathbf{1}_{\Omega_{n}}\lVert V_{n+1}(X_{n})\rVert^{2}\\ &=(1-\gamma_{n}c^{2})D_{n}\mathbf{1}_{\Omega_{n}}+\gamma_{n}\xi_{n+1}\mathbf{1}_{\Omega_{n}}+\frac{L\gamma_{n}^{2}}{2}\mathbf{1}_{\Omega_{n}}\lVert V_{n+1}(X_{n})\rVert^{2},\\ &\leq{D}_{1}\prod_{k=1}^{n}(1-\gamma_{k}c^{2})+\sum_{k=1}^{n}\left(\prod_{j=k}^{n}(1-\gamma_{j}c^{2})\right)\gamma_{k}\xi_{k+1}\mathbf{1}_{\Omega_{n}}\\ &\quad+\frac{L}{2}\sum_{k=1}^{n}\left(\prod_{j=k}^{n}(1-\gamma_{j}c^{2})\right)\gamma_{k}^{2}\lVert V_{k+1}(X_{k})\rVert^{2}\mathbf{1}_{\Omega_{n}}\\ &\leq{D}_{1}\prod_{k=1}^{n}(1-\gamma_{k}c^{2})+\sum_{k=1}^{n}\left(\prod_{j=k}^{n}(1-\gamma_{j}c^{2})\right)\gamma_{k}\xi_{k+1}\mathbf{1}_{\Omega_{n}}\\ &\quad+\frac{L}{2}\sum_{k=1}^{n}\gamma_{k}^{2}\lVert V_{k+1}(X_{k})\rVert^{2}\mathbf{1}_{\Omega_{n}},\end{split}$ (32)